# Victoria, BC, Canada

### v1

@@ Abstract
Mixed and multi methods research is often used in software engineering, but researchers outside of the social or human sciences often lack experience when using these designs. This paper provides guidelines and advice on how to design mixed and multi method research, and to encourage the intentional, rigourous, and innovative use of mixed methods in software engineering. It also presents key characteristics of core mixed method research designs. Through a number of fictitious but recognizable software engineering research scenarios and personas of prototypical researchers, we showcase how to choose suitable designs and consider the inevitable tradeoffs any design choice leads to. We furnish the paper with recommended best practices and several antipatterns that illustrate what to avoid in mixed and multi method research.
@@
Mixed methods Multi methods Research methods Methodology Guidelines
@@ 1 Introduction
Software engineering (SE) researchers use a variety of research methods, often mixing and matching them. Despite the widespread prevalence of mixing methodsbryman_integrating_2006 ; creswell_best_2011 ; poth_innovation_2018 ; ladner_mixed_2019 ; poth_sage_2023 , there is little guidance in SE on how to do so systematically, in a way that is well motivated, carefully designed, robustly executed, and ideally, leading to additional benefits not otherwise achievable from the use of a single research method alone. Mixing methods is particularly pertinent in today’s complex socio-technical SE contexts where multiple methods can bring insights on the nuanced nature of the problem or solution being studied.
Our paper is aimed at researchers unfamiliar with mixed methods. We share key concepts and provides guidance in how to mix and match methods, presenting different characteristics of mixed methods and some common designs. Inspired by our previous paper on selecting empirical methods for software engineering research easterbrook_selecting_2008 , this paper provides a number of illustrative fictitious research scenarios that capture different research aims and call for very different research design choices. These scenarios help to guide design choice but also emphasize the tradeoffs faced in making mixed method study design decisions. As well as showcasing design choices for a variety of scenarios, the paper also calls out best practices and a number of antipatterns to be avoided.
@@
The timing of this paper is important, as budding SE researchers today face a complex world where software is used to address complex problems with wide-scale societal impact. There is also a strong drive for innovation in the face of rapidly changing technologies and practices, all occurring in socio-technical contexts storey_who_2020 ; hoda_socio-technical_2021 . The increasing interest in studying the adoption of generative AI technologies in everyday SE practice is but one latest example of the constantly changing SE research landscape bano2023LLM . Choosing which mixed methods to use and knowing how to use them effectively is challenging, and is often overlooked in SE research and education curricula and thus, experience is lacking in how to use these methodological designs.
In this paper, we include practical illustrative examples in the form of four fictional SE research personas and scenarios, representing common mixed method contexts, designs, and outcomes. These four personas and their research scenarios are used to exemplify and motivate the use of mixed and multiple method research designs throughout the paper. The personas we introduce include:
  • Ali, a postdoc studying how the addition of security experts to development teams impacts developer motivations and behaviours;
  • Sam, an early stage Ph.D. student studying the impact of a new code review tool at a small company;
  • Vicki, an experienced industry researcher studying the adoption of generative AI at their company; and
  • Zara, a late stage Ph.D. student evaluating an automated tool they designed.
@@
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed and multi methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel insights, procedural rigor and ethical research. In Section 4, we discuss some notable characteristics of mixed method designs, as well as common mixed and multi method research designs. In Section 5, we present the four personas of budding new researchers in software engineering and detail the challenging research decisions they face. We describe the methodological choices they make and how they followed the four core principles mentioned in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research before concluding the paper in Section 7. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
@@ 2 A Brief Overview of Mixed Methods Research
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that MMR gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, “The formative period of mixed methods began in the 1950s and continued up until the 1980s. This period saw the initial interest in using more than one method in a study” but these authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until much later maxwell_expanding_2016 . In the social sciences field, Campbell and Fisk’s paper in 1959campbell_convergent_1959 is recognized as the first adoption of multiple research methods for validation purposes. They also defined this design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et.al webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s, discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. johnson_toward_2007 who coined mixed methods research as the third paradigm, coexisting with quaN and quaL studies. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods, i.e., which describe how to structure and design a MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
@@
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark poth_sage_2023 when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of a MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences become central.
Figure 1: Interconnection of the core components for MMR: collecting, analyzing and interpreting (1) qualitative and (2) quantitative data to address a research question; (3) integrating the data (merging, connecting, and building) in a research design and set of procedures; and (4) drawing meta-inferences that add new insight beyond the quantitative and qualitative data. Adapted from Creswell and Plano Clark poth_sage_2023 .
@@
3 Principles to Guide Mixed and Multi Method Research
Figure 2: Principles for Mixed and Multi Method Research Designs
Building on Poth poth_innovation_2018 , we define mixed and multi methods research (MMMR) [^1] Expanding the MMR abbreviation to include an additional ‘M’ for multi. as, A research approach where multiple procedures are used to collect, analyze, and integrate data to address a research problem. For mixed methods, both qualitative and quantitative data are integrated to produce novel insights.
The difference between mixed methods and multi methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi methods can be applied with any combination including quaN-quaN or quaL-quaL.
Poth poth_innovation_2018 also presents four core guiding principles for MMMR that we have adapted slightly to be applicable to software engineering (see Figure 2). Two of the principles, methodological rationale and novel insights, relate to understanding why we should consider using a mixed or multiple methods research design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying mixed and multi methods research. We delve into these four principles below.
@@ 3.1 Methodological Rationale
Principle of Methodological Rationale: MMMR study designs explain how the mixed and multiple research design and data procedures are relevant for addressing a given research problem. An MMMR study description should answer the question why did we select a particular MMMR research design and set of procedures?
There are several conceptual reasons for using mixed or multiple methods [^2] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_social_2012 , Creswell creswell_designing_2018 and McGrath mcgrath_methodology_1995 . The rationale for the methodology may be one or more of the following:
@@
  • Credibility: another procedure may be used to increase the integrity and truth of the findings from a former procedure bryman_social_2012 . Multiple or mixed methods can increase credibility by bringing multi-dimensional philosophical and theoretical lenses to the findings mertens_mixed_2013 .
@@
3.2 Novel Insights
Principle of Novel Insights: MMMR studies generate new insights through the integration of the procedures, which in turn addresses the reason(s) for the use of multiple or mixed methods. A well-executed MMMR study answers the question what did we gain from using MMMR?
Succinctly, an MMMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method:
@@ 3.3 Procedural Rigor
Principle of Procedural Rigor: Mixed and multi method research studies enhance the reliability and validity of the findings throughout the study. A well-executed MMMR study answers the question how well did we conduct the study?
Rigor in any research study is always a critical concern, and it is often a difficult concept for students to know how to assess. Rigor is particularly challenging for mixed and multi methods. MMMR studies should first adhere to the rigor of the individual studies, but there are also some concerns that apply to the “mixing” aspect. For this principle, we are guided by a mixed methods rigor appraisal tool called MMAT (Mixed Methods Appraisal Tool)  oliveira_mixed_2021 . The following are posed as questions to be considered when self-reflecting on the rigor of a mixed design:
@@
  • Are the findings from the mixed or multiple methods integrated? One must ask what are the similarities and divergencies in the findings (the output of the methods), asking does the integration add to more than the sum of the multiple parts? Any contradictions or divergences should be explained or discussed.
  • Are the different methods used rigorously conducted? MMMR studies should adhere to the rigor of the individual studies used in the mixed or multi study (referring to the appropriate quality criteria for each study).
@@ 3.4 Ethical Research
Principle of Ethical Research: MMMR studies manifest the application of responsible research, with respect for people (e.g., participants, concerned parties, beneficiaries) and their welfare (e.g., respect for their confidentiality, privacy, time, emotions, cultural sensitivity, and needs), and that of the environment (e.g., organisations, communities of practice, society, and the planet at large, aiming for sustainability). A good MMMR study answers the question how responsibly did we do it?
Ethical research is a timeless concept. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, guidelines, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of any research work.
An ethical approach to research starts by carefully considering the why and not being driven by the why not. This applies to the use of any procedure, but is even more pertinent to consider with mixed or multi methods because of the additional risks these designs bring in terms of privacy and confidentiality when mixing and integrating different types of data.
@@
  • Privacy and confidentiality: Deep and revealing conversations with research participants lend richness and nuance to the findings. Similarly, access to sensitive data for research purposes, such as team chats, emails, bug reports, and financial reports, is a privilege. What are our plans for protecting participant and data privacy and confidentiality when gathering, storing, and processing these assets, and when presenting findings based on them?
  • Respect and cultural sensitivity: Many studies involve people from multiple locations and cultural backgrounds. Have we considered these cultural sensitivities when designing our protocols? Do we need to customise our protocols to make better sense to different participant groups or be more culturally appropriate in some communities? Respect extends to a participant’s time and privacy, e.g., how will we ensure the privacy of traceable information leading to individual identification in open-source repositories or in end-user reviews of software apps and products? Some types of mining of publicly available data can be considered unintended and an unapproved use, especially where the chain of evidence can lead to identification and loss of privacy.
  • Safety and welfare: A cornerstone of ethical research is to ensure the safety and welfare of the people and environments touched by the research. What are the physical, virtual, or psychological risks we may be exposing our participants and ourselves to? What are the environmental implications of our research? Do we have the requisite training to conduct risk assessments and to ensure the safety of our participants, ourselves, and the environment?
@@
  • Special concerns about MMMR: While many of the ethical concerns above may equally apply to a single method research study, there are some that may be raised specifically due to the use of MMMR. Typically, MMMR studies collect and analyze a variety of data from multiple sources. While this is a strength of MMMR studies, it is also a potential threat. Because of the multiple angles at which an MMMR study approaches a phenomenon, mixing at the data level can place participant privacy and confidentiality under threat. For example, if recruiting survey participants for follow-up interviews from a survey cohort, it may be feasible to trace interviews back to survey responses using demographic data. Similarly, interview responses can be traced back to email, chat, or code commit logs, jeopardising the privacy and confidentiality of certain participants or data sources. Thus, special care needs to be taken when using MMMR.
4 A Landscape of Mixed and Multi Methods Design

### v2

@@ Abstract
Mixed methods research is often used in software engineering, but researchers outside of the social or human sciences often lack experience when using these designs. This paper provides guidelines and advice on how to design mixed method research, and to encourage the intentional, rigorous, and innovative use of mixed methods in software engineering. It also presents key characteristics of core mixed method research designs. Through a number of fictitious but recognizable software engineering research scenarios, we showcase how to choose suitable designs and consider the inevitable trade-offs any design choice leads to. We describe several antipatterns that illustrate what to avoid in mixed method research, and when mixed method research should be considered over other approaches.
@@
Mixed methods Research methods Methodology Guidelines
@@ 1 Introduction
Software engineering (SE) researchers use a variety of research methods, often mixing and matching them. Despite the widespread prevalence of mixing methodsbryman_integrating_2006 ; creswell_best_2011 ; poth_innovation_2018 ; ladner_mixed_2019 ; poth_sage_2023 , there is little guidance in SE on how to do so systematically, in a way that is well motivated, carefully designed, robustly executed, and leads to additional benefits not otherwise achievable from the use of a single research method alone. Mixing methods is particularly pertinent in today’s complex socio-technical SE contexts storey_who_2020 ; hoda2024Book where quantitative and qualitative methods, when combined, can bring insights on the nuanced nature of the problem or solution being studied.
Although mixed methods are known in SE, there is not enough debate on the methodological issues, challenges and trends that set the current scenario of mixed methods research in SE. Di Penta and Tamburri Dipenta discuss how mixed methods in empirical SE combine quantitative and qualitative research, and how qualitative methods can enhance quantitative studies and vice versa. Critically, there is no clear guidelines explaining mixed methods research for SE. To the best of our knowledge, our paper is the first to provide guidelines on mixed methods research designs in an SE context that researchers can adopt when designing studies that foresee mixing qualitative (quaL) and quantitative (quaN) methods.
Our paper is aimed at researchers unfamiliar with mixed methods (the use of both quaN and quaL methods) and those keen to improve their understanding and practice. We share key concepts and provide guidance on how to mix and match methods, presenting different characteristics of mixed methods and some common designs.
@@
The timing of this paper is important, as SE researchers today face a complex world where software is used to address complex problems with wide-scale societal impact. There is also a strong drive for innovation in the face of rapidly changing technologies and practices, all occurring in socio-technical contexts storey_who_2020 ; hoda_socio-technical_2022 . The increasing interest in studying the adoption of generative AI technologies in everyday SE practice is but one latest example of the constantly changing SE research landscape bano2023LLM . Choosing which mixed methods to use and knowing how to use them effectively is challenging, and is often overlooked in SE research and education curricula and thus, experience is lacking in how to use these methodological designs.
Inspired by a previous paper, one of us co-authored on selecting empirical methods for SE research easterbrook_selecting_2008 , this paper includes practical illustrative examples in the form of four fictional SE researchers and research scenarios, representing common mixed method contexts, designs, and outcomes. These four research scenarios are used to exemplify and motivate the use of mixed method research designs throughout the paper. While not exhaustive, these scenarios not only help to guide design choice but also emphasize the trade-offs faced in making some common mixed method study design decisions. As well as showcasing design choices for a variety of scenarios, the paper also calls out best practices and a number of antipatterns that should be avoided.
@@
The rest of this paper is organized as follows. In Section 2, we present a brief background of the most prominent and relevant influences in mixed methods research design over the decades. In Section 3, we present four core principles that underlie mixed method research: methodological rationale, novel integrated insights, procedural rigor and ethical research. In Section 4, we discuss some notable characteristics of mixed method designs, as well as common mixed methods research designs. In Section 5, we present four scenarios of research studies that fictional researchers pursue. We describe the methodological choices these characters make and how they followed the four core principles mentioned in Section 3. In Section 6, we use the concept of antipatterns to present what not to do in mixed method research. In Section 7, we delve in to the importance of integration in mixed methods research, discuss mixed methods research designs with other research designs (such as multi-method, case studies and secondary studies), and share a call to action before concluding the paper in Section 8. This paper includes an Appendix which contains a glossary of terms we use throughout the paper for ease of reference.
@@ 2 A Brief Overview of Mixed Methods Research
It is only in the mid-20th century, with roots in disciplines such as sociology and anthropology, that mixed methods research (MMR) gained traction with the formalization of approaches by authors including Creswell and Plano Clark creswell_designing_2018 . According to these authors, “The formative period of mixed methods began in the 1950s and continued up until the 1980s. This period saw the initial interest in using more than one method in a study” but these authors did not explicitly identify these approaches as mixed methods. Furthermore, the different typologies of mixed method designs did not emerge until the 1990s maxwell_expanding_2016 ; Greene1989 .
In the social sciences field, Campbell and Fisk’s paper in 1959 campbell_convergent_1959 is recognized as the first adoption of multiple research methods for validation purposes. They also defined this design as multi-operationalism since the authors propose a multitrait, multimethod matrix that intersects quaL-quaN methods with a specific set of traits as a means for enforcing validity of research conclusions. In the mid 1960s, Webb et al. webb_unobtrusive_1966 introduced the concept of triangulation, followed by Denzin in the 1970s denzin_research_1978 who detailed how to triangulate using both within-methods triangulation and between-methods triangulation, emphasizing the latter as a more powerful means to overcome bias and as a way to pragmatically mix methods.
Although the combination of quaL-quaN methods has been used in practice for many years, the first evidence of a more formal use of mixed methods is identifiable only in the late 1980s, discussed by Guest and Fleming guest_mixed_2015 and Johnson et al. johnson_toward_2007 who coined mixed methods research as the third paradigm, coexisting with quaL and quaN studies. The common denominator in these early conceptualizations of mixed methods is that MMR combines quaL-quaN elements. It is only recently that researchers have turned their attention to proposing typologies of mixed methods, i.e., which describe how to structure and design a MMR by combining quaL and quaN elements according to some criteria as stated by Teddlie and Tashakkori teddlie_foundations_2009 .
@@
With the increasing complexity of MMR designs, the problem of discussing and visually presenting complex designs to readers became apparent. Consequently, as noted by Creswell and Plano Clark (poth_sage_2023, , Chapter 2) when they revisited MMR designs twenty years after their important chapter was first published in 2003 tashakkori_handbook_2003 , the most recent studies decreased their focus on the quaN and quaL components and increased attention towards the integration processes across design types. To reflect this change about the central elements of the MMR designs, Figure 1 represents the simplification of the interconnection of the core components of a MMR design. From this perspective, treating the data beyond quaL and quaN analysis and gaining insight through integration and meta-inferences become central.
Figure 1: Interconnection of the core components for MMR: collecting, analyzing and interpreting (1) qualitative and (2) quantitative data to address a research question; (3) integrating the data (merging, connecting, and building) in a research design and set of procedures; and (4) drawing meta-inferences that add new insight beyond the quaL and quaN data. Adapted from Creswell and Plano Clark (poth_sage_2023, , Chapter 2).
@@
3 Principles to Guide Mixed Method Research
Building on Poth poth_innovation_2018 , we define mixed methods research (MMR) as, a research approach where multiple procedures are used to collect, analyze, and integrate data to address a research problem. For mixed methods, both qualitative and quantitative data are integrated to produce novel insights.
Poth poth_innovation_2018 also presents four core guiding principles for MMR that we have adapted slightly to be applicable to SE (see Figure 2). Two of the principles, methodological rationale and novel integrated insights, relate to understanding why we should consider using a MMR design. The other two, procedural rigor and ethical research, relate to understanding how we should go about designing and applying MMR. While some aspects of these principles apply to single method applications and can be considered good research practice in general, others are particularly relevant to MMR and generally more complex in that context. We delve into these four principles below.
Figure 2: Principles for Mixed Method Research Designs, as inspired by Poth poth_innovation_2018 .
@@ 3.1 Methodological Rationale
Principle of Methodological Rationale: Mixed method research study designs should explain how the mixed research design and data procedures are relevant for addressing a given research problem. By adhering to this principle, an MMR study description can address the question why did we select a particular MMR research design and set of procedures?
There are several conceptual reasons for using mixed methods [^1] The reasons we list are inspired by or earlier proposed by Ladner ladner_mixed_2019 , Bryman bryman_social_2012 , Creswell creswell_designing_2018 and McGrath mcgrath_methodology_1995 . The rationale for the methodology may be one or more of the following:
@@
  • Credibility: another procedure may be used to increase the integrity and truth of the findings from a former procedure bryman_social_2012 . Mixed methods can increase credibility by bringing multi-dimensional philosophical and theoretical lenses to the findings mertens_mixed_2013 .
@@
3.2 Novel Integrated Insights
Principle of Novel Integrated Insights: Mixed methods research studies should generate new insights through the integration of the procedures, which in turn addresses the reason(s) for the use of mixed methods. An MMR study report should clarify the novelty of the study findings and explain how methods and/or data were integrated. By adhering to this principle, an MMR study report can address the question what did we gain from using MMR?
Succinctly, an MMR study should deliver on its promise of additional benefits to what could have been afforded by using a single method:
@@ 3.3 Procedural Rigor
Principle of Procedural Rigor: Mixed method studies should apply the study procedures rigorously and explain them effectively. By adhering to this principle, an MMR study report can address the question how well did we conduct the study?
Rigor in any research study is always a critical concern, and it is often a difficult concept for students to know how to assess. Rigor is particularly challenging for mixed methods. MMR studies should first adhere to the rigor of the individual studies, but there are also some concerns that apply to the “mixing” aspect. For this principle, we are guided by a mixed methods rigor appraisal tool called MMAT (Mixed Methods Appraisal Tool) oliveira_mixed_2021 . The following are posed as questions to be considered when self-reflecting on the rigor of a mixed design:
@@
  • Are the findings from the mixed methods integrated? One must ask what are the similarities and divergencies in the findings (the output of the methods), asking does the integration add to more than the sum of the parts? Any contradictions or divergences should be explained or discussed.
  • Are the different methods used rigorously conducted? MMR studies should adhere to the rigor of the individual studies used in the mixed study (referring to the appropriate quality criteria for each study [^4] See the ACM Empirical Standards sigsoft ).
@@ 3.4 Ethical Research
Principle of Ethical Research: Mixed methods research studies should manifest the application of responsible research, with respect for people (e.g., participants, concerned parties, beneficiaries) and their welfare (e.g., respect for their confidentiality, privacy, time, emotions, cultural sensitivity, and needs), and that of the environment (e.g., organisations, communities of practice, society, and the planet at large, aiming for sustainability). By adhering to this principle, an MMR study report can address the question how responsibly did we do it?
Ethical research is a timeless concept. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, guidelines, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of research.
An ethical approach to research starts by carefully considering the why and not being driven by the why not. This applies to the use of any procedure, but is even more pertinent to consider with mixed methods because of the additional risks these designs bring in terms of privacy and confidentiality when mixing and integrating different types of data.
@@
  • Privacy and confidentiality: Deep and revealing conversations with research participants lend richness and nuance to the findings. Similarly, access to sensitive data for research purposes, such as team chats, emails, bug reports, and financial reports, is a privilege. Typically, MMR studies collect and analyze a variety of data from multiple sources. While this is a strength of MMR studies, it is also a potential threat. Because of the multiple angles at which an MMR study approaches a phenomenon, mixing at the data level can place participant privacy and confidentiality under threat. For example, if recruiting survey participants for follow-up interviews from a survey cohort, it may be feasible to trace interviews back to survey responses using demographic data. Similarly, interview responses can be traced back to email, chat, or code commit logs, jeopardising the privacy and confidentiality of certain participants or data sources. What are our plans for protecting participant and data privacy and confidentiality when gathering, storing, and processing these assets, and when presenting findings based on them?
  • Respect and cultural sensitivity: Many studies involve people from multiple locations and cultural backgrounds. Have we considered these cultural sensitivities when designing our protocols? Do we need to customise our protocols to make better sense to different participant groups or be more culturally appropriate in some communities? Similarly, have we considered what data sovereignty may apply to special groups such as Indigenous peoples [^6] https://static1.squarespace.com/static/5b3043afb40b9d20411f3512/t/63ed934fe861fa061ebb9202/ 1676514134724/Communique-Indigenous-Data-Sovereignty-Summit.pdf?
  • Safety and welfare: A cornerstone of ethical research is to ensure the safety and welfare of the people and environments touched by the research. For example, asking participants to recall adverse experiences or imagine challenging scenarios may cause stress. What are the physical, virtual, or psychological risks we may be exposing our participants and ourselves to? What are the environmental implications of our research? Do we have the requisite training to conduct risk assessments and to ensure the physical, psychological, and mental safety and wellbeing of our participants, ourselves, and the environment?
4 A Landscape of Mixed Methods Design